OpenStudy (anonymous):

Solve M = 2x + 3y for y.

OpenStudy (anonymous):

Use Algebra.

Parth (parthkohli):

^

OpenStudy (anonymous):

Particularly this: \[A=B+C \Rightarrow A-B=C\]

Parth (parthkohli):

Isolate y.

OpenStudy (anonymous):

And this: \[\alpha \beta=\gamma \Rightarrow \alpha=\frac{\gamma}{\beta}\]

OpenStudy (amistre64):

Deld - if you were to attempt this, how would you try to start solving it? right or wrong, what is your ideas